OpenStudy (anonymous):

sqrt(2x-1)=5-sqrt(x-1) solve for x

OpenStudy (anonymous):

2x-1=25-10sqrt(x-1)+(x-1)

OpenStudy (anonymous):

2x-1-24-x=-10sqrt(x+1) x-25=-10sqrt(x+1)

OpenStudy (anonymous):

x^2-50x+625=100(x+1) I trust you can solve from here

OpenStudy (anonymous):

how did you get 10 please

OpenStudy (anonymous):

5-sqrt(something)*(5-sqrt[something]) 25*-5sqrt[something]-5sqrt[something]+something 25-10sqrt[something]+something